Grade 11Mechanics

If x=3t^2-8t+5, find the velocity of the particle when the displacement of the particle is zero

Kashish
9 Years agoGrade 11

1 Answer

V Kartik
ApprovedApproved Tutor Answer9 Years ago
3t^2-8t+5=0==>t=5/3 or t=1 v= dx/dt=>v= (6t-8)/1 or 3(6t-8)/5=>v= -2 ms^-1 or 6/5 ms^-1Understood